OpenStudy (anonymous):

Solve. Please show work. x+1/5=2/3.

OpenStudy (turingtest):

LCD if 1/5 and 2/3=15 x+(1/5)(3/3)=2/3(5/5) x+3/15=10/15 x=7/15